Anisha is 25 years younger than her uncle.In 3 years time,her uncle will be six times her age.find her age
Answer by Timnewman(323)

let her uncle's age be x,
Anisher's age =x-25
Tn 3yrs time,
Uncle age=x+3
Anisher's age=x-22
Then,
x+3=6(x-22)
x+3=6x-132
6x-x=132+3
5x=135
x=27
Uncle's age is 27yrs,
Anisher's age is 27-25=2yrs